OpenStudy (anonymous):

if f(x)=x^2+3x+5, what is f(3+h)

OpenStudy (amistre64):

plug it in and simplify to your hearts desire .... whats the problem you are having?

OpenStudy (anonymous):

I do not know which part of the equation I am supposed to plug it into. Would it be into the x-value?

OpenStudy (amistre64):

compare: f(x) to f(3+h) what has changed?

OpenStudy (amistre64):

yes, the x parts are being defined by the notation, think of it as f(x=3+h)

OpenStudy (amistre64):

\[f(x)=x^2+3x+5\] \[f(3+h)=(3+h)^2+3(3+h)+5\]

OpenStudy (anonymous):

So, it would be stretched out into (3+h)^2+3(3+h)+5? Would simplifying that give me the answer?

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

Ok, so would it be h^2+9h+23?

OpenStudy (amistre64):

maybe, lets chk 3^2 + h^2 + 6h + 3(3) + 3h + 5 looks good to me

OpenStudy (anonymous):

Thank you so much!

OpenStudy (amistre64):

youre welcome